A race car accelerates on a straight road from rest to speed of 180 km/hr in 25s . Assuming uniform acceleration of the car through out find the distance covered in this time .Using only graphically method.

Grade:11

2 Answers

KARTIK GUPTA
6 years ago
Given,
u = 0 ,v= 180 km/hr =50 m/s , t = 25 seconds
We know,
v = u + at
50 = a×25
a = 2 m/sec2
Distance covered by the body,
S = ut + at2 /2
  =  0+ 2×625/2
  = 625 m = 0.625 km
 
Chiranjeet
5 years ago
Given: u=0;v=180km/h=50m/s;t=25s.
From 1st Equation of Motion,
v=u+at
50=0+a×25
a=2m/s^2.
From 3rd Equation of Motion,
s=ut+ 1/2at^2
  =0×25+1/2×2×(25)^2
  =625m=0.625km